What is the specific heat capacity of W if 145.8 J were required to heat 46.10 g of it from 25.0°C to 49.32 °C?
1
Expert's answer
2020-12-16T07:14:41-0500

C = "\frac{Q}{MΔT}"


C = specific heat

Q = energy = 145.8 J

ΔT (change in temperature) = 49.32 ⁰C - 25.0 ⁰C = 24.32 ⁰C

M (mass) = 46.10 g


C = "\frac{145.8 J}{46.10×10^-³(24.32⁰C)}"

C = 130.0448 J/Kg•⁰C